# SNOOPI: Supercharged One-step Diffusion Distillation with Proper Guidance

### v2

@@ 4.1 Motivation
While the VSD framework offers a solid foundation for model distillation, its training process often experiences instability. For example, while successfully working on Stable Diffusion 2.1 (SDv2.1) [38], our re-implementation of SwiftBrush and SwiftBrushv2 (SB) [33, 7] fail in distilling other model backbones such as Stable Diffusion 1.5 (SDv1.5) [38] and PixArt-$\alpha$ [6] when using a fixed CFG scale in the VSD framework. Their unstable results are substantiated by quantitative evaluation shown in Fig. 2.
Figure 3: Qualitative comparison of different methods for integrating negative prompts into the generation process using a one-step generator. As shown, simply applying CFG to a one-step generator is equivalent to blending (i.e., adding) two images together, resulting in an unusable output. Alternatively, negating the negative embedding from the positive embedding shows minimal impact on the final image. In contrast, NASA is the first method to successfully steer the generation away from negative attributes in a one-step generator, while also producing high-quality results.
To address this instability, previous studies have proposed techniques such as using image datasets as regularizers, increasing the LoRA teacher update frequency, and adding adversarial loss terms [54, 55]. However, these methods have two main drawbacks: (1) they rely on image supervision, and (2) they significantly increase memory usage and training time. In pursuit of image-free distillation, we instead investigate the existing SB framework to identify the root cause of the instability.
We begin by examining the impact of the CFG scale in multi-step diffusion models, specifically SDv2.1. Our analysis of the MS COCO 2014 benchmark [24], with results shown in Tab. 1, reveals that model performance shifts considerably as the guidance scale changes. Typically, users set the guidance scale based on their specific goal, whether to prioritize image fidelity or text alignment. However, this approach is less desirable in training, as we want the model to be versatile across different scenarios. Given this, we seek to make the teacher model adaptable across various scenarios without compromising overall performance. As shown in Tab. 1, varying the guidance scale randomly allows us to achieve competitive results across all metrics while maintaining flexibility across different guidance levels.
$\kappa$ FID $\downarrow$ CLIP $\uparrow$ Precision $\uparrow$ Recall $\uparrow$
2 10.77 0.31 0.54 0.53
3 9.60 0.32 0.59 0.50
4 10.35 0.32 0.60 0.47
5 11.53 0.33 0.61 0.44
$\mathcal{U}(2,5)$ 9.86 0.32 0.60 0.48
Table 1: Results on the zero-shot MS COCO-2014 30K benchmark [24] of SDv2.1 [38] when using PNDM Scheduler with different CFG scale $\kappa$.
Teacher LoRA Teacher FID $\downarrow$
10.72
10.81
9.51
9.03
Table 2: Ablation studies on Random CFG in VSD loss with SDv2.1 backbone on 2K prompts from MS COCO-2014 [24].
4.2 Method
Based on our findings, we aim to broaden the guidance distribution provided by the teacher in the SB distillation framework by introducing Proper Guidance - SwiftBrush (PG-SB) which applies a randomized guidance scale during teacher inference. Specifically, rather than fixing $\kappa=C$, we sample its value from a uniform distribution $\boldsymbol{\mathcal{U}(\kappa_{\text{min}},\kappa_{\text{max}})}$ where $\kappa_{\text{min}}$ and $\kappa_{\text{max}}$ are predefined constants. In the SB distillation pipeline, however, CFG is applied in two places: once to the frozen teacher and once to the LoRA teacher. As shown in Tab. 2, applying this technique only to the LoRA teacher improves model performance. However, applying it solely to the frozen teacher destabilizes the distillation process and eventually harms the student performance. The gradient of our PG-SB’s VSD loss is defined by:
$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}^{\text{PG-SB}}=&\mathbb{E}_{t,\mathbf{y},\mathbf{z},\boldsymbol{\kappa\sim\mathcal{U}(\kappa_{\text{min}},\kappa_{\text{max}})}}\bigg{[}\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\phi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg{]}.\end{split}$ (7)
@@ 4.2 Method
We hypothesize that broadening the output distribution of the LoRA teacher enables it to better align with the student’s output, mainly since the student may not initially perform well at one function evaluation or one NFE generation. Interestingly, once the LoRA teacher’s output distribution is expanded, the student can also adapt to the broader range of outputs from the frozen teacher when using a randomized guidance scale, achieving the best results when varied guidance scales are applied to both teachers.
5 Negative-Away Steer Attention
5.1 Motivation
Negative prompting is an essential technique for improving control in text-to-image and text-to-3D synthesis. This method helps exclude unwanted elements in generated images and mitigate the Janus problem in 3D generation [20, 31, 1]. However, no existing approach supports negative prompts in single-step or few-step diffusion models.
@@ 5.1 Motivation
A widely method for implementing negative prompting relies on the CFG mechanism. However, applying CFG directly to one-step diffusion models presents significant challenges. CFG was designed for multi-step models, where the output is iteratively refined, allowing for gradual adjustments using both negative and unconditional outputs. In contrast, one-step models generate images in a single pass without the iterative denoising process, making them inherently less flexible. Directly applying CFG to these models leads to an undesirable blending of images, resulting in unnatural artifacts, as shown in Fig. 3. This highlights a crucial limitation: the CFG mechanism, though effective in multi-step models, is not directly translatable to one-step models without significant modifications.
An alternative approach manipulates text embeddings by subtracting the negative embedding from the positive one to steer the model away from undesired attributes. However, as shown in Fig. 3, this method has minimal effect on the generated images, highlighting the need for a more robust solution to negative prompting in one-step models.
5.2 Method
@@
Instead of applying CFG in the input space (text embedding) or output space (synthesized image) as in multi-step diffusion, we propose moving CFG to the intermediate feature space of the diffusion model. We hypothesize that cross-attention layer features are particularly well-suited for this approach. Specifically, cross-attention layers in these models capture semantic connections between image regions and text features [13], making them ideal for controlling feature emphasis. By modulating these cross-attention maps, we can filter unwanted features by subtracting attention responses to negative prompts from those to positive prompts. This filtering occurs in the feature space, meaning that undesired elements are excluded before the image is generated, reducing the chance of blending artifacts.
@@
Given the latent’s features at current layer $l$, $\mathbf{Z}_{l}$, and positive text features $\mathbf{c}_{p}$ extracted from CLIP text encoder, we define the query, key, and value matrices as $\mathbf{Q}_{l}=\mathbf{Z}_{l}\mathbf{W}_{q,l}$, $\mathbf{K}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{k,l}$, and $\mathbf{V}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{v,l}$, where $\mathbf{W}_{q,l}$, $\mathbf{W}_{k,l}$, and $\mathbf{W}_{v,l}$ are projection matrices at current layer $l$. For negative text features $\mathbf{c}_{n}$, we use the same projection matrices, defining the key and value matrices as $\mathbf{K}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{k,l}$ and $\mathbf{V}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{v,l}$.
@@
where $\alpha$ is the scale factor to control the degree of negative feature removal. The overall pipeline of the NASA mechanism is illustrated in Fig. 4.
@@ 6.1 Experimental Setup
Metrics and benchmark. In diffusion model evaluation, Fréchet Inception Distance (FID) [14] on the zero-shot MS-COCO2014 benchmark has traditionally been the standard metric. However, recent studies [37, 5, 54] suggest that FID does not always correlate with actual image quality and adopt human-focused alignment metrics like Human Preference Score v2 (HPSv2) [49] for benchmarking. Our experiments also follow this conclusion, so we choose HPSv2 as the primary metric for evaluating the models. Additionally, we use CLIP scores to assess text alignment and precision/recall [22] to evaluate general image quality and mode coverage. For completeness, we still report FID results. Except for HPSv2, all metrics are computed following the protocols in [18, 7]: generating 30K images using MS-COCO2014 prompts, resizing them to $256\times 256$, and comparing with center-cropped images from the full dataset.
@@
Implementation details. We base our method on SBv2 [7] with our proposed modifications on CFG scale $\kappa$. We use total of 3.3M prompts from JourneyDB [47] and a subset of LAION [43]. All our training is conducted on four NVIDIA A100 40GB GPUs with the total batch size of 64. The learning rates are set to $1e^{-6}$ for the student model and $1e^{-3}$ for the LoRA teacher, using the AdamW optimizer [27]. We utilize three backbones SDv1.5 [38], SDv2.1 [38] and PixArt-$\alpha$ [6] as the frozen teachers and LoRA teachers are initialized with rank $r=64$ and scaling $\gamma=128$. The teacher models’ CFG scales are set as follows: $\kappa_{\text{min}}=0.5,\kappa_{\text{max}}=4$ for SDv1.5 and SDv2.1 and $\kappa_{\text{min}}=0.5,\kappa_{\text{max}}=3$ for PixArt-$\alpha$. Further details can be found in the Appendix.
6.2 Evaluation of PG-SB
Tab. 4 and Tab. 5 present a thorough quantitative comparison between our approach and prior distilled text-to-image diffusion models. The results consistently show the superiority of our model in HPSv2 scores, with the highest average score across all tested backbones. This reinforces the robustness of our method, as discussed in Sec. 4.2 and demonstrated in Fig. 2, while also revealing a key insight. By gaining a deep understanding of the underlying framework, we made a simple yet impactful change with minimal overhead, leading to significant performance improvements. This suggests that integrating complex loss functions may not always be necessary for achieving state-of-the-art results. Remarkably, our method achieves these outcomes without requiring image-level supervision, restoring SBv2 as an image-free distillation approach.
NASA SDXL-LCM SDXL-DMD2 SDXL-DMD2 PG-SB
4 steps 4 steps 1 step 1 step
43% 27% 25% 38%
97% 100% 100% 92%
Table 3: Comparison of success rate of unwanted feature removal in generated images between models with and without NASA.
Figure 5: Effect of different scale values (0.0 to 1.0) in NASA with SDXL-DMD2 model, illustrating the progressive influence on visual details and composition.
Figure 6: Qualitative results of NASA. Each row displays images generated with a positive (green) and negative (red) prompt pair. Models using NASA effectively exclude unwanted features, aligning outputs with desired traits.
Method Anime Photo Concept Art Paintings Average
Stable Diffusion 1.5-based backbone
SDv1.5 [38] 26.51 27.19 26.06 26.12 26.47
InstaFlow-0.9B [26] 26.10 26.62 25.92 25.95 26.15
DMD2 [54] 26.39 27.00 25.80 25.83 26.26
PG-SB 27.18 27.58 26.69 26.62 27.02
PG-SB + NASA 27.19 27.59 26.71 26.63 27.03
Stable Diffusion 2.1-based backbone
SDv2.1 [38] 27.48 26.89 26.86 27.46 27.17
SB [33] 26.91 27.21 26.32 26.37 26.70
SBv2 [7] 27.25 27.62 26.86 26.77 27.13
PG-SB 27.56 27.84 26.97 27.03 27.35
PG-SB + NASA 27.71 27.99 27.14 27.27 27.53
PixArt-$\alpha$-based backbone
PixArt-$\alpha$ [6] 29.62 29.17 28.79 28.69 29.07
YOSO [30] 28.79 28.09 28.57 28.55 28.50
DMD [55] 29.31 28.67 28.46 28.41 28.71
PG-SB 32.19 29.09 30.39 29.69 30.34
PG-SB + NASA 32.56 29.55 31.24 30.96 31.08
Table 4: HPSv2 comparisons between our method and previous works. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
6.3 Evaluation of NASA
To assess the effectiveness of NASA, we conduct experiments with our model and other one-step and few-step models, such as LCM [29] and DMD2 [54]. We design a small set of positive and negative prompts and generate 100 images per model, both with and without negative prompts (further details are provided in the Appendix). We then calculate the percentage of images that successfully exclude the feature specified by the negative prompt. For example, given the positive prompt “A photo of a person” and the negative prompt “male”, we measure the percentage of generated images that do not depict a male person. The average results across all prompt sets are summarized in Tab. 3. Our findings show that NASA is compatible with both one-step and few-step models, enabling precise control over unwanted features in generated images. Note that the difference in accuracy between PG-SB and other models is primarily attributed to the disparity in backbone size. The SDXL-based architecture is well-known for its superior text alignment, which stems from its larger generator backbone and the use of an ensemble of text encoders. Additionally, Fig. 5 and Fig. 6 present qualitative results of our method. Images generated with a conditional negative prompt retain key semantic features such as pose and background, while filtering out unwanted attributes specified by the negative prompt. Furthermore, when using negative prompts like “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark”, our method improves image quality, as shown in Tab. 4 and Tab. 5. Notably, integrating NASA with PG-SB on the PixArt-$\alpha$-based backbone achieves a record-breaking HPSv2 score of 31.08, surpassing all other one-step diffusion models baselines.
Method #Params NFEs FID $\downarrow$ CLIP $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ Image-free?
Stable Diffusion 1.5-based backbone
SDv1.5 ($\textit{cfg}=3$) [38] 0.9B 25 8.78 0.30 0.59 0.53
UFOGen [51] 0.9B 1 12.78 - - -
InstaFlow-0.9B [26] 0.9B 1 13.33 0.30 0.53 0.45
DMD [55] 0.9B 1 11.49 0.32 - -
DMD2 [54] 0.9B 1 8.29 0.30 0.59 0.52
PG-SB 0.9B 1 10.08 0.31 0.57 0.47
PG-SB + NASA 0.9B 1 9.94 0.31 0.57 0.48
Stable Diffusion 2.1-based backbone
SDv2.1 ($\textit{cfg}=3$) [38] 0.9B 25 9.60 0.32 0.59 0.50
SB [33] 0.9B 1 15.46 0.30 0.47 0.46
SBv2 [7] 0.9B 1 8.14 0.32 0.57 0.52
PG-SB 0.9B 1 8.91 0.34 0.57 0.49
PG-SB + NASA 0.9B 1 8.83 0.34 0.57 0.50
PixArt-$\alpha$-based backbone
PixArt-$\alpha$ ($\textit{cfg}=4.5$) [6] 0.6B 20 26.85 0.32 0.52 0.23
YOSO [30] 0.6B 1 26.04 0.30 0.46 0.30
DMD [55] 0.6B 1 30.22 0.32 0.55 0.18
PG-SB 0.6B 1 22.03 0.33 0.52 0.24
PG-SB + NASA 0.6B 1 21.98 0.33 0.52 0.25
Table 5: Quantitative comparisons between our method and others on zero-shot MS COCO-2014 benchmark. For multi-step SD models, we report each with the CFG scale that returns the best FID. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints. ‘-’ denotes unreported results.

### v3

@@ 4.1 Motivation
Figure 2: The inherent single-pass nature of one-step diffusion models renders them incompatible with direct CFG, resulting in artifact-laden outputs.
Positive prompt is not enough. While vision generative models like diffusion models excel at incorporating positive attributes, they fundamentally struggle with understanding negation. Consider requesting “A portrait of a friendly monster, not scary” (Fig. 1), the model typically fails to exclude the specified negative attribute “scary”. Negative prompting addresses this limitation by creating a separate, dedicated input for exclusions – completely removing negation terms from the positive prompt and processing them independently. This clear separation allows the model to accurately distinguish what to avoid generating, providing significantly better control over the final output.
Community Need. The growing use of negative prompting in the generative AI community further highlights its importance. Platforms like Civit AI [10] and various open-source repositories [11, 2, 29] have shown that using both types of prompts together results in more refined, higher-quality model outputs that align more closely with user specifications. Users can clearly specify both desired and unwanted attributes in a prompt pair, which is particularly useful when working with models that generate complex or highly detailed content. Experimentally, Table 1 and Figure 1 confirm that using negative prompts effectively enhances image quality of the model.
Critical Gap. Despite the established need and widespread adoption of negative prompts in multi-step diffusion models and the rapidly growing prominence of one-step approaches, a significant limitation persists: no existing one-step methods effectively support negative prompting. Our proposed method fills this gap by providing a simple, efficient, and training-free mechanism that integrates negative prompt into one-step diffusion models to (1) improve controllability and (2) enhance the quality of the output image.
@@
4.2 Challenges
@@
A widely used method for implementing negative prompting is the CFG mechanism. However, applying CFG directly to one-step diffusion models presents significant challenges. CFG was designed for multi-step models, where the output is iteratively refined, allowing for gradual adjustments using both negative and unconditional outputs. In contrast, one-step models generate images in a single pass without the iterative denoising process, making them inherently unsuitable for this method. Direct application of CFG to these models leads to undesirable image blending, resulting in unnatural artifacts, as shown in Fig. 2. This highlights a crucial limitation: the CFG mechanism, though effective in multi-step models, is not directly translatable to one-step models without significant modifications.
5 Proposed Method
Our proposed methodology is detailed in Sec. 5.1 for the core mechanism. We present two primary implementation strategies for NASA: a training-free approach (NASA–I) and a distillation-based approach (NASA–T), which are presented in Sec. 5.2 and Sec. 5.3, respectively.
5.1 NASA: Negative-Away Steer Attention
In contrast to CFG implementations that operate within the output space (synthesized image), as commonly practiced in multi-step diffusion models, we propose a novel approach that strategically relocates the guidance mechanism to the intermediate representation space of the diffusion model.
@@ 5.1 NASA: Negative-Away Steer Attention
Specifically, cross-attention layers in these models capture semantic connections between image and text features [17], making them ideal for controlling the feature alignment. We hypothesize that cross-attention layer features are particularly well-suited for this approach. We selectively manipulate these cross-attention maps to attenuate attention responses associated with negative prompts relative to those elicited by positive prompts, thereby filtering out unwanted semantic attributes.
This representation-space guidance strategy offers a crucial advantage: it enables the exclusion of undesirable elements prior to the final image synthesis stage. This removal mechanism significantly mitigates the risk of generating blending artifacts, which are common byproducts of image-space manipulation techniques.
Our representation-space guidance also extends to architectures like FLUX [25], which replace traditional cross-attention with joint self-attention blocks. [47] show that these blocks still embed distinct image-text cross-attention operations. This allows NASA to be seamlessly applied, proving its robustness across different backbones.
@@
The overall pipeline of the NASA mechanism is illustrated in Fig. 3. Given the latent feature representation at a given layer $l$, denoted as $\mathbf{Z}_{l}$, and positive text features $\mathbf{c}_{p}$ extracted from CLIP text encoder, we define the query, key, and value matrices as $\mathbf{Q}_{l}=\mathbf{Z}_{l}\mathbf{W}_{q,l}$, $\mathbf{K}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{k,l}$, and $\mathbf{V}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{v,l}$, where $\mathbf{W}_{q,l}$, $\mathbf{W}_{k,l}$, and $\mathbf{W}_{v,l}$ are projection matrices at current layer $l$. Analogously, for negative text features $\mathbf{c}_{n}$, we use the same projection matrices to define the corresponding key and value matrices for negative prompt conditioning: $\mathbf{K}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{k,l}$ and $\mathbf{V}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{v,l}$.
Method FLUX.1-schnell SDXL-LCM SDXL-DMD2
4 steps 1 step 4 steps 1 step 4 steps 1 step
None 23% 44% 43% - 27% 25%
CFG 30% 0% 14% - 25% 0%
NASA–I 100% 99% 97% - 100% 100%
Table 2: Comparison of success rate of unwanted feature removal in generated images. Note that SDXL-LCM [31] does not support one-step inference.
@@
where $\alpha$ is the scale factor to control the degree of negative feature removal.
5.2 NASA–I: Training-free Setting
Applicable to One-/Few-Step Models. To assess the effectiveness of NASA–I, we carry out experiments with pretrained one-step and few-step models, such as LCM [31], DMD2 [51] and FLUX.1-schnell [25]. We design a small set of positive and negative prompts and generate 100 images per model, both with and without negative prompts. We then calculate the percentage of images that successfully exclude the feature specified by the negative prompt. For example, given the positive prompt “A photo of a person” and the negative prompt “male”, we measure the percentage of generated images that do not depict a male person. The average results across all prompt sets are summarized in Tab. 2. Our findings show that while traditional CFG is ineffective for one-step and few-step models, NASA–I provides precise control over unwanted features in generated images. Additionally, this capability is also qualitatively demonstrated in Fig. 1, Fig. 4 and Fig. 5.
Efficiency Considerations. As described in Sec. 5.1, we reuse the projection matrices of the positive prompt for the negative prompt, ensuring that no additional model parameters are introduced and keeping the model size unchanged. Furthermore, applying NASA-I to all one-step and few-step models introduces only a minor computational overhead, increasing FLOPs by just 1.89 %. This represents a significant efficiency improvement over CFG, which doubles the computational cost due to the need for separate forward passes for positive and negative prompts. Consequently, NASA–I enables negative prompt guidance in one-step diffusion models without introducing substantial latency, making it a practical solution for real-time applications.
5.3 NASA–T: Distillation Setting
Given their effectiveness in improving image quality during inference, their potential utility in model distillation is worth exploring. In this section, we introduce a method that integrates the negative prompt into one-step diffusion distillation. While our approach is applicable to various distillation methods, we select SwiftBrush (SB) [34, 13], an image-free method, as the baseline for improvement due to its efficiency. As described in Eq. 6, SB updates the student model using the gradient of the VSD loss. To incorporate the negative prompt $\mathbf{y}^{-}$, we modify the existing formulation by replacing $\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)$ (see Eq. 4) with $\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\mathbf{y}^{-},\kappa)$ (see Eq. 5). In this way, the teacher model provides explicit guidance to the student model, helping it learn to steer away from undesirable features. Moreover, we incorporate the NASA mechanism into the student model, enabling it to process negative prompts during training. This allows the student model to better recognize and suppress unnatural artifacts, leading to improved learning and enhanced image quality. Note that with the LoRA teacher model $\hat{\boldsymbol{\epsilon}}_{\pi}$, we can choose whether to incorporate negative prompts. Empirically, we found that not using them yields better results. Detailed results and discussions are provided in Sec. 6.3.
@@ 6 Experiments
Dataset NegOpt HPSv2
Method CLIP+ $\uparrow$ CLIP- $\downarrow$ Anime $\uparrow$ Photo $\uparrow$ CA $\uparrow$ Paintings $\uparrow$ Average $\uparrow$
PixArt-$\alpha$-based backbone
PixArt-$\alpha$ [9] [Teacher] 0.35 0.05 29.62 29.17 28.79 28.69 29.07
YOSO [32] 0.36 0.08 28.75 28.06 28.52 28.57 28.48
+ NASA–I 0.36 0.06 28.74 28.05 28.56 28.60 28.49 (+0.01)
DMD [52] 0.35 0.08 29.31 28.67 28.46 28.41 28.71
+ $\text{CFG}=1.5$ 0.34 0.09 30.02 27.07 28.36 28.07 28.38 (-0.33)
+ $\text{CFG}=2.5$ 0.31 0.13 26.74 23.86 25.13 24.66 25.10 (-3.61)
+ NASA–I 0.35 0.05 29.33 28.71 28.49 28.53 28.77 (+0.06)
SBv2 0.36 0.09 32.19 29.09 30.39 29.69 30.34
+ NASA–I 0.36 0.06 32.60 29.58 31.09 30.65 30.98 (+0.64)
+ NASA–T 0.35 0.08 32.33 29.26 30.75 30.10 30.61 (+0.27)
+ NASA–T + $\text{CFG}=1.5$ 0.34 0.10 29.47 26.50 28.22 27.68 27.97 (-2.37)
+ NASA–T + NASA–I 0.35 0.06 32.66 29.67 31.31 30.71 31.09 (+0.75)
+ NASA–T + NASA–I ($\alpha=1$) 0.35 0.05 32.65 29.65 31.45 31.06 31.21 (+0.87)
Stable Diffusion 1.5-based backbone
SDv1.5 [40] [Teacher] 0.31 0.06 26.51 27.19 26.06 26.12 26.47
InstaFlow-0.9B [28] 0.33 0.11 26.10 26.62 25.92 25.95 26.15
+ NASA–I 0.33 0.10 26.15 26.68 25.93 25.98 26.19 (+0.04)
DMD2 [51] 0.31 0.10 26.39 27.00 25.80 25.82 26.25
+ NASA–I 0.31 0.09 26.39 27.02 25.80 25.83 26.26 (+0.01)
SBv2 0.32 0.09 27.18 27.58 26.69 26.62 27.02
+ NASA–I 0.32 0.08 27.19 27.59 26.71 26.63 27.03 (+0.01)
+ NASA–T 0.35 0.10 27.64 27.94 27.19 27.03 27.45 (+0.43)
+ NASA–T + NASA–I 0.35 0.09 27.65 27.97 27.18 27.04 27.46 (+0.44)
Stable Diffusion 2.1-based backbone
SDv2.1 [40] [Teacher] 0.33 0.06 27.48 26.89 26.86 27.46 27.17
SBv2 [13] 0.34 0.10 27.25 27.62 26.86 26.77 27.13
+ NASA–I 0.35 0.08 27.44 27.84 26.91 27.02 27.30 (+0.17)
SBv2 0.37 0.10 27.56 27.84 26.97 27.03 27.35
+ NASA–I 0.37 0.08 27.70 28.00 27.12 27.22 27.51 (+0.16)
+ NASA–T 0.36 0.09 28.00 28.65 27.44 27.26 27.84 (+0.49)
+ NASA–T + NASA–I 0.36 0.07 28.04 28.68 27.51 27.50 27.93 (+0.58)
Table 3: Quantitative comparisons between our method and previous works. NASA–I refers to our training-free method introduced in Sec. 5.2, while NASA–T denotes our distillation-based approach described in Sec. 5.3. SBv2 refers to our reimplementation of SBv2 with a randomly sampled CFG scale $\kappa$.
Figure 4: Effect of different scale values (0.0 to 1.0) in NASA-I with SDXL-DMD2 [51] for both one-step and few-step settings, illustrating the progressive influence on visual details and composition.
Figure 5: Qualitative results of NASA-I in FLUX.1-schnell [25].
Figure 6: Qualitative results of NASA-I for text-to-video generation in CausVid [53].
Method VBench-Long benchmark
Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$
None 61.98 67.12
NASA–I 63.33 67.36
Table 4: CausVid quality metrics.
@@ 6.1 Experimental Setup
Metrics and Benchmark. In diffusion model evaluation, Fréchet Inception Distance (FID) [18] on the zero-shot MS-COCO2014 [27] benchmark has been the standard metric. However, recent studies [39, 8, 51] suggest that FID in this benchmark does not always correlate with actual image quality and adopt human-focused alignment metrics like Human Preference Score v2 (HPSv2) [49] for benchmarking. Our experiments also follow this conclusion, so we choose HPSv2 as the primary metric for evaluating the models. Moreover, due to the absence of negative prompts in the MS-COCO2014 benchmark, we utilize the NegOpt dataset from [35], which contains pairs of positive and negative prompts. From this dataset, we randomly sample 30K prompt pairs to construct a subset for evaluation. Text alignment is then assessed using CLIP scores, where CLIP+ evaluates alignment with the positive prompt and CLIP- evaluates alignment with the negative prompt.
Training Datasets. Following SBv2 [13], we use total of 3.3M prompts from JourneyDB [45] and a subset of LAION [43] [^1] All datasets were downloaded and evaluated at MovianAI.. For negative prompts in the distillation setting (Sec. 5.3), we construct a dataset of 10K negative prompts derived from commonly used terms like “bad photo, duplicate, low resolution, low quality”.
@@
Implementation Details. We base our method on SBv2 [13] with our proposed modifications in Sec. 5.3. We utilize three backbones SDv1.5 [40], SDv2.1 [40] and PixArt-$\alpha$ [9] as the frozen teachers and LoRA teachers are initialized with rank $r=64$ and scaling $\gamma=128$. The learning rates are set to $1e^{-6}$ for the student model and $1e^{-3}$ for the LoRA teacher, using the AdamW optimizer [30]. All our training is conducted on four NVIDIA A100 40GB GPUs with the total batch size of 64 for the SD models and 32 for PixArt-$\alpha$. Moreover, [6] found that instead of using a fixed CFG scale $\kappa$ during training, sampling $\kappa$ from a uniform distribution within a given range improves training robustness. Based on this insight, we set the teacher models’ CFG scales as follows: $\kappa\sim\mathcal{U}(0.5,4)$ for SDv1.5 and SDv2.1, and $\kappa\sim\mathcal{U}(0.5,3)$ for PixArt-$\alpha$. For NASA–T, in each iteration, we uniformly sample the scale $\alpha\sim\mathcal{U}(0,1)$ to help the student model adapt to different conditions and enhance generalization. For NASA–I, unless explicitly specified, we set $\alpha=0.1$ for SDv1.5-based models, $\alpha=0.2$ for SDv2.1-based models, and $\alpha=0.5$ for PixArt-$\alpha$-based models. Further details can be found in the Appendix.
NASA
Teacher
LoRA Teacher
HPSv2 Avg. $\uparrow$ 30.34 29.84 29.80 30.61 29.89
Table 5: Ablation studies on the integration of negative prompts in model distillation using the PixArt-$\alpha$ backbone.
Figure 7: Ablation study on the NASA scale using our distilled model based on the PixArt-$\alpha$ backbone.
6.2 Main Results
Quantitative Results. Table 3 presents a comparison between our approach and prior distilled text-to-image diffusion models. Using our reimplementation of SBv2 [13] as the baseline, NASA–T consistently enhances image quality across all tested backbones. Compared to other methods, our model achieves the highest HPSv2 scores on average, demonstrating its clear superiority in image generation quality. This suggests that integrating complex loss functions like DMD2 [51] may not always be necessary for achieving state-of-the-art results. Moreover, while applying CFG directly degrades one-step model outputs, NASA–I further enhances image quality across all models when using negative prompts like “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark”. Notably, our method on the PixArt-$\alpha$-based backbone achieves a record-breaking HPSv2 score of 31.21, surpassing all other one-step diffusion models baselines. Regarding the NegOpt dataset, with the support of NASA–I, all models demonstrate a reduction in the CLIP- score while maintaining the CLIP+ score. This indicates our method effectively eliminates unwanted features in generated outputs while preserving desired traits.
Qualitative Results. As discussed in Sec. 5.2, NASA–I is a versatile, training-free method for both one-step and few-step diffusion models. Figure 4 and Figure 5 present qualitative results of integrating NASA–I into the SDXL-DMD2 [51] and FLUX.1-schnell [25] models, respectively. When using a conditional negative prompt, the generated images retain key semantic features while effectively suppressing unwanted attributes specified by the negative prompt.
Application to Text-to-Video Models. To demonstrate the versatility of our approach, we extended NASA-I to a few-step video diffusion model, CausVid [53]. As shown in Fig. 6, our method effectively suppresses unwanted visual attributes while preserving key semantic content and temporal consistency. Quantitative evaluation on the VBench-Long [22] benchmark further validates this. Table 4 shows that using NASA–I with common negative prompts like “low quality, bad anatomy, blurry, distortion, ugly, low resolution, unclear details” improves both the aesthetic and imaging quality scores of the generated videos. This highlights NASA–I’s potential as a general-purpose tool for enhancing guided generation across different modalities.
6.3 Ablation Studies
Table 5 presents an ablation on integrating negative prompts during model distillation with the PixArt-$\alpha$ [9] backbone. The baseline is our reimplementation of SBv2 with a randomly sampled CFG scale $\kappa$. Simply applying negative prompts only in the teacher and/or the LoRA teacher models slightly degrades the HPSv2 score, as the student model is not explicitly trained on this negative conditioning. In contrast, incorporating NASA with negative prompts during training allows the student model to better capture and utilize this conditioning, leading to improved performance. Furthermore, we find that integrating negative prompts solely into the teacher model yields better results than applying them to both the teacher and LoRA teacher models.
Using our PixArt-$\alpha$-based distilled model as the base, we varied the NASA scale $\alpha$ and evaluated the results with HPSv2 on the benchmark dataset. Figure 7 demonstrates that NASA–I generally enhances image synthesis, resulting in outputs that better align with human preferences. However, if the NASA scale $\alpha$ is set too high, the negative attributes may overly dominate the positive attributes, adversely affecting the final image quality.